OpenStudy (anonymous):

Consider the function below. f(x) = 3 - 3x - x^2 Evaluate the difference quotient for the given function. Simplify your answer. f(2+h)-f(2)/h

jimthompson5910 (jim_thompson5910):

\[\large f(x) = 3 - 3x - x^2\] \[\large f(2) = 3 - 3(2) - (2)^2\] \[\large f(2) = 3 - 3(2) - 4\] \[\large f(2) = -7\] --------------------------------------------------------- \[\large f(x) = 3 - 3x - x^2\] \[\large f(2+h) = 3 - 3(2+h) - (2+h)^2\] \[\large f(2+h) = 3 - 3(2+h) - (4+4h+h^2)\] \[\large f(2+h) = 3 - 6-3h - 4-4h-h^2\] \[\large f(2+h) = -7-7h-h^2\] -------------------------------------------------------- The key things to draw out from the work above is that \[\large f(2) = -7\] and \[\large f(2+h) = -7-7h-h^2\] \[\large \frac{f(2+h)-f(2)}{h}\] \[\large \frac{-7-7h-h^2-(-7)}{h}\] \[\large \frac{-7-7h-h^2+7}{h}\] \[\large \frac{-7h-h^2}{h}\] \[\large \frac{h(-7-h)}{h}\] \[\large \frac{\cancel{h}(-7-h)}{\cancel{h}}\] \[\large -7-h\] \[\large -h-7\] So \[\large \frac{f(2+h)-f(2)}{h}=-h-7\]